The steady state temperature of certain medium is given by theta = e powerΒ 2x - 3 y. Find the linear approxmiation at (0,0)


1
Expert's answer
2022-03-10T03:58:56-0500
"\theta=e^{2x-3y}=e^{2x}*e^{-3y}=(1+(2x)+O(x^2))*(1-3y+O(y^2))=\\
=1+2x-3y+O(x^2)+O(x*y)+O(y^2)"

linear approximation at (0,0) is equal to 1+2x-3y